A painter leans a ladder against one wall of a house. The ladder is 16ft long. The base of the ladder is 10ft from the house. How high does the ladder reach on the wall of the house?
Answer by RAY100(1637)

a quick sketch helps clarify.
This is a right triangle problem, ground is perpendicular to bldg, one leg is 10, another is height, and hypotenuse is length of ladder, 16.
use pythagorous to resolve, c^2=a^2 +b^2
16^2=10^2 +h^2
16^2-10^2=h^2
h^2=156
h=12.5
check, 12.5^2+10^2=16^2 ok